# Analysis of Discrete Modern Hopfield Networks in Open Quantum System

### v1

@@ I Introduction
The modern Hopfield network, devised by Krotov and Hopfield, generalizes the energy function of the conventional Hopfield network [7], significantly increasing its storage capacity [7, 8, 9, 10]. Because of its generality, the modern Hopfield network contains other prominent learning models, such as the transformer [9, 11] used in natural languages, and diffusion models [12, 13] used in computer vision, as special cases. When discretized, a modern Hopfield network functions equivalently to a neural network with one hidden layer [7].
@@
Another approach involves embedding the network into a dissipative quantum spin system, governed by the quantum master equation [21, 22, 23, 24, 25]. Owing to stochastic time evolution, these models can avoid stacking at the local minima of the energy function, similar to annealing algorithms. In the time evolution of this type of quantum Hopfield network, a transverse field serves as a quantum effect, including non-classical behavior of solutions [21] and altering storage capacity[25]. Generalizing this quantum framework to modern Hopfield networks could enhance efficacy and applicability; however, existing studies are limited to extensions of conventional Hopfield networks.
@@
In our model, the energy function takes an integer power function, aligning with the Hopfield network when it is quadratic. Notably, our model exhibits qualitatively distinct behavior when the exponent exceeds two and when the exponent is two (i.e., when the model is an open quantum Hopfield network [21]). When the exponent exceeds two, there always is a stable steady state and the phase diagram has a region described by limit cycle where the origin is always stable.
The reminder of this paper is organized as follows. Sec. II introduces our proposed model. Sec. III presents the findings on the phase diagram, solution behavior, and storage capacity under varying quantum and temperature parameters. Sec. IV summarizes our results and discusses future issues.
@@ II.1 Discrete Hopfield Network
A discrete Hopfield network is defined on a graph consisting of $N$ nodes and edges connecting each node pair. This model is a recurrent (the output of the node at time $t$ becomes the input at time $t+1$) and asynchronous (update nodes one at a time, and their updates do not affect the other nodes) model, with a fully connected network. Each node takes the value of ±$\pm 1$, which we interpret as classical 1/2-spins.
@@
The time evolution in (1) decreases the system’s energy. This energy minimization is expected to lead the network to retrieve a memory pattern as the energy eventually minimizes.
@@ II.3 Open Quantum Generalization of Discrete Modern Hopfield Network
$\Delta E_{k}:=\sum_{j_{1}\ldots j_{x-1}}W_{k,j_{1}\ldots j_{x-1}}\sigma_{j_{1}}^{Z}\cdots\sigma_{j_{x-1}}^{Z},$ (12)
is the change of “energy” under flipping the $k$-th spin. Selecting $x=2$ in Eq. 4 recovers the quantum Hopfield network introduced in Ref. [21]. The quantum effect of this model are governed by the transverse magnetic field Hamiltonian:
@@ II.4 Time evolution of averaged spins
By inserting Eqs. (11), (13) into the above time evolution and ignoring the terms of order $\mathcal{O}(N^{-(x-1)})$, we obtain that [21]
@@
By multiplying both sides of Eqs. (17), (18) by $\frac{1}{N}\sum_{i}\xi_{i}^{\mu}$, we obtain
@@
[…]
$\left\langle\frac{1}{N}\sum_{i=1}^{N}\xi_{i}^{\mu}\tanh\left(\beta\sum_{\mu}\xi^{\mu}\left(\langle M_{Z}^{\mu}\rangle\right)^{x-1}\right)\right\rangle$
@@
[…]
@@ III Results
This study analyzes the phase diagram of the proposed model. The phase diagram of an open quantum Hopfield network is examined in Ref. [21], and its storage capacity is analyzed in Ref. [25]. We extend these results analytically and numerically, focusing primarily on the case where $x=4$ and $p=1$.
@@ III.1 Phase Diagram
This section presents the phase diagram of our model for $p=1$ (one stored memory). Initially, we analyze the number of fixed points in the time evolution of $M_{Z}$ (24) and $M_{Y}$ (25) as follows:
@@
Additionally, we have $M_{Y}=-4\Omega M_{Z}$. Notably, the origin $(M_{Z},M_{Y})=(0,0)$ is always a fixed points. The boundary at which the number of fixed points changes is determined when the left-hand side (LHS) and right-hand side (RHS) of Eq. 28 are tangential:
@@
The parameters $(T,\Omega)$ that satisfies these two equations provide the desired boundary.
@@ III.1.1 $x=2$
This defines the boundary where the number of the fixed points varies. When $\beta<\beta_{c}$, the origin is the only fixed; however, when $\beta>\beta_{c}$, three fixed points exist.
We also numerically examine the region with a LC, as shown in Fig. 1, where the area above the black line corresponds to a region with a single solution, while the area below has two or more solutions. However, after examining the stability of each solution, we identify a region below the black line where the ferromagnetic (FM) solution is unstable. Therefore, we perform a simulation and obtain the green and blue regions, as shown in Fig. 1. The green region in Fig. 1 represents the LC phase, while the blue region is the FM+LC phase. Furthermore, the boundary obtained by calculating the stability of the FM solutions in Ref. [21] do not coincide with the boundary between the blue and green regions in Fig. 1.
@@
Concequently, the stability of the PM solution is given by the following conditions [21]:
@@
    $\{\beta<\beta_{c}\}\cap\{\beta>3/2\}\cap\{\beta>4\Omega+1/2\}$: Since the eigenvalues are positive real value, the solution is unstable.
@@
Condition 1 and 2 indicate that the PM solution is unstable and is classified as the FM phase. Condition 3 indicates instability with spiraling behavior and is classified as the LC phase. Condition 4 and 5 indicate stability and is classified as the PM phase.
Figure 1: The phase diagram of the quantum modern discrete Hopfield network ($x=2$) in $(T,\Omega)$ plane. The black line indicates the boundary (31) where the number of solutions changes. The blue+green region signify the area where LCs are detected from an initial value away from the origin ($(M_{Z},M_{Y})=(3,-3)$). The green region is the area where LCs are detected from an initial value around the origin ($(M_{Z},M_{Y})=(0.05,-0.05)$). Alternatively, the blue region is where the initial values outside the LC converges to it, while the initial value inside the LC and close to the origin converges to the stable fixed points, indicating the FM + LC area. The numbering from $(1)$ to $(4)$ corresponds to Table 1.
@@ III.1.2 $x=4$
which determines the boundary at which the number of the fixed points changes. When $\beta>\beta_{c}$, there are five fixed points instead of three owing to the non-trivial power inside $\tanh(\cdot)$.
@@
Figure 2: The phase diagram of the quantum modern discrete Hopfield network ($x=4$) in $(T,\Omega)$ plane. The black line indicates the boundary Eq. 36, where the number of solutions changes. The blue region is the area where the simulation suggests the presence of a LC when the initial value is $(M_{Z},M_{Y})=(3,-3)$. Only one stable attractor in the PM phase (above the curve) and the FM phase has two or more stable attractors (below the curve). The label LC + PM indicates the region where a LC appears and the origin remains stable.
@@
This transformation allows us to obtain the stability around the solution, classifying it similar to the case for $x=2$. However, this boundary changes slightly from the existing boundary.
In addition to the fixed points, we numerically identify some parameter regions with a nontrivial LC encircling the origin (blue region in Fig. 2). However, unlike in the open quantum Hopfield network, the LC consistently coexists with at least one stable fixed point (origin), and the shape of the region in the phase space differs qualitatively from that of the quantum Hopfield network.
Figure 3: Enlarged view of the region where the LC appears. $(1)$ to $(4)$ represent the number of Table 1. $(2)$ refers the blue region above the black line and $(3)$ refers the narrow blue region below the black line. The region $(3)$ contains two different phases, depending on whether the fixed points with $M_{Z}>0$ are stable (FM+LC) or unstable (PM+LC).
Number $x=2$ [21] $x=4$ (this paper)
$(1)$ FM FM
$(2)$ LC PM + LC
$(3)$ FM + LC FM + LC    /PM + LC
$(4)$ PM PM
Table 1: Comparison of the phase at $x=2$ and $x=4$.
@@ III.2.1 Paramagnetic (PM) phase
The PM phase corresponds to the region where every initial state reaches the origin, which is the only stable fixed point.
III.2.2 Paramagnetic + Limit Cycle (PM+LC) phase
@@
The PM+LC phase corresponds to the region where the only stable fixed point is the origin; however, there is an additional LC around it. In a quantum Hopfield network, the LC phase does not coexist with the PM phase; thus, any initial $(M_{Z},M_{Y})$ converges to the LC. However, in our model, the LC always coexists with other stable fixed points. In the PM+LC phase, the initial point converges to the origin or the LC depending on its proximity to the origin.
III.2.3 Ferromagnetic (FM) phase
@@
The region with stable fixed points, such that $M_{Z}>0$, but without LC is termed as the FM phase. Unlike the FM phase of a quantum Hopfield network [21], which only has stable fixed points with $M_{Z}>0$, our model exhibits an FM phase with an additional stable fixed point at the origin. Consequently, if the initial $M_{Z}$ is sufficiently small, the magnetization $M_{Z}$ diminishes, similar to the PM phase. In addition, the FM region in Fig. 2 is smaller than that in a quantum Hopfield network [21].
III.2.4 Ferromagnetic + Limit Cycle (FM+LC) phase
@@
The FM+LC phase corresponds to the region where FM and LC coexist. In this region, LC encapsulate all stable fixed points. When the initial point lies within the LC, it converges to one of the fixed points or the LC, depending on the closeness to these points. Conversely, any point outside the LC converges to the LC.
III.2.5 The behavior of the fixed points for larger $x$
@@
We also analyze the time evolution for $x$ values exceeding $4$ Fig. 6. In a discrete modern Hopfield network, memory capacity increases nonlinearly with $N$ as $x$ increases [7]. The time evolutions of the averaged spins varies with increasing $x$ in an open quantum discrete Hopfield network. The point of change is that the range of $M_{Z}$ that converges to the origin expands as the value of $x$ increases.
@@
The open quantum discrete modern Hopfield network is described by Eq. 26 and Eq. 27, and the behavior of these functions changes only with $x$ via $\tanh(\beta(M_{Z})^{x-1})$. The function $\tanh(\beta(M_{Z})^{x-1})$ has no extreme value at $x=2$, but has an extreme value $M_{Z}=0$ when $x>2$, indicating that the origin is always stable. Consequently, the behavior of the model changes qualitatively between $x=2$ and $x>2$, but not significantly for $x>2$.
@@
Figure 4: The time evolution of the averaged spin for each parameter $T,\Omega$ in the FM and PM phases. The red dots are the stable fixed points. (a), (b), and (c) represent the time evolution in the FM phase, and (d) and (e) represent the time evolution in the PM phase.
@@
Figure 5: The time evolution of the averaged spin for each parameter $T,\Omega$ in the phase in which the LC appears.(f), (g), and (h) represent the time evolution of the averaged spin in the PM+LC Phase, and (i) represents the time evolution of the averaged spin in the FM+LC phase.
@@ III.3 Storage Capacity
Two methods analyze the storage capacity of Hopfield networks. The first calculates the maximum number of patterns that can be stored without unacceptable errors [1, 28]. The second employs statistical physics analysis [5, 28]. Notably, the storage capacity of Hopfield network is $0.14N$.
The storage capacity of a discrete modern Hopfield network is $\alpha_{x}N^{x-1}$ where $\alpha_{x}$ is a numerical constant that depends on the threshold of error [7]. Moreover, models with more generalized energy functions and time evolution can achieve an exponential storage capacity [8, 9]. The storage capacity of a quantum Hopfield network at a given temperature $T$ and quantum drive $\Omega$ is also known [25]. The maximal capacity $\alpha_{c}$ of our model can be calculated using the method described in Ref. [25]. While the classical and quantum limits coincide, further exploration of quantum capacity remains a future challenge.

### v2

@@ I Introduction
The modern Hopfield network, devised by Krotov and Hopfield, generalizes the energy function of the conventional Hopfield network [7], significantly increasing its storage capacity [7, 8, 9, 10]. Due to its generality, the modern Hopfield network contains other prominent learning models, such as the transformer [9, 11] used in natural languages, and diffusion models [12, 13] used in computer vision, as special cases. When discretized, a modern Hopfield network functions equivalently to a neural network with one hidden layer [7].
@@
Another approach involves embedding the network into a dissipative quantum spin system, governed by the quantum master equation [21, 22, 23, 24, 25]. Due to stochastic time evolution, these models can avoid stacking at the local minima of the energy function, similar to annealing algorithms. In the time evolution of this type of quantum Hopfield network, a transverse field serves as a quantum effect, including non-classical behavior of solutions [21] and altering storage capacity[25]. Generalizing this quantum framework to modern Hopfield networks could enhance efficacy and applicability; however, existing studies are limited to extensions of conventional Hopfield networks.
@@
In our model, the energy function takes an integer power function, aligning with the Hopfield network when it is quadratic. Notably, our model exhibits qualitatively distinct behavior when the exponent exceeds two and when the exponent is two (i.e., when the model is an open quantum Hopfield network [21]). When the exponent exceeds two, there is always a stable steady state, and the phase diagram has a region described by a limit cycle (LC) where the origin is always stable.
The remainder of this paper is organized as follows. Section II introduces our proposed model. Section III presents the findings on the phase diagram, solution behavior, and storage capacity under varying quantum and temperature parameters. Section IV summarizes our results and discusses future issues.
@@ II.1 Discrete Hopfield Network
A discrete Hopfield network is defined on a graph consisting of $N$ nodes and edges connecting each node pair. This model is a recurrent (the output of the node at time $t$ becomes the input at time $t+1$) and asynchronous (update nodes one at a time, and their updates do not affect the other nodes) model, with a fully connected network. Each node takes the value of $\pm 1$, which we interpret as classical 1/2-spins.
@@
The time evolution in Eq. (1) decreases the system’s energy. This energy minimization is expected to lead the network to retrieve a memory pattern as the energy eventually minimizes.
@@ II.3 Open quantum generalization of discrete modern Hopfield network
$\Delta E_{k}:=\sum_{j_{1}\ldots j_{x-1}}W_{k,j_{1}\ldots j_{x-1}}\sigma_{j_{1}}^{Z}\cdots\sigma_{j_{x-1}}^{Z}$ (12)
is the change of “energy” under flipping the $k$-th spin. Selecting $x=2$ in Eq. 4 recovers the quantum Hopfield network introduced in Ref. [21]. The quantum effects of this model are governed by the transverse magnetic field Hamiltonian:
@@ II.4 Time evolution of averaged spins
By inserting Eqs. (11) and (13) into this time evolution and ignoring the terms of order $\mathcal{O}(N^{-(x-1)})$, we obtain that [21]
@@
By multiplying both sides of Eqs. (17) and (18) by $\frac{1}{N}\sum_{i}\xi_{i}^{\mu}$, we obtain
@@
We assume an unbiased independent and identically distributed random memory pattern ($\xi^{\mu}_{i}=\pm 1$ with a probability $1/2$). Consider the average and the evolution of $\langle M_{a}^{\mu}\rangle$, where $\langle\cdot\rangle$ denotes the expected value for randomly stored memory patterns. Because the graph is the complete graph, the mean field approximation $\langle M_{a}^{\mu}M_{b}^{\nu}\rangle\approx\langle M_{a}^{\mu}\rangle\langle M_{b}^{\nu}\rangle$ becomes exact with the limit of $N\to\infty$ [2]. Similar to Ref. [21], we assumed the following self-averaging hypothesis:
$\left\langle\frac{1}{N}\sum_{i=1}^{N}\xi_{i}^{\mu}\tanh\left(\beta\sum_{\mu}\xi^{\mu}_{i}\left(\langle M_{Z}^{\mu}\rangle\right)^{x-1}\right)\right\rangle$
@@
For simplicity, we omit $\langle\cdot\rangle$ and denote $M_{a}^{\mu}\equiv\langle M_{a}^{\mu}\rangle$.
@@ III Results
This study analyzes the phase diagram of the proposed model. The phase diagram of an open quantum Hopfield network is examined in Ref. [21], and its storage capacity is analyzed in Ref. [25]. We extend these results analytically and numerically, focusing primarily on the case where x=4x=4111We choose $x=4$ to preserve the spin-inversion symmetry of the energy function. Odd values such as $x=3$ break this symmetry and make it difficult to analytically derive the relationship between $\beta$ and $\beta_{c}$. and $p=1$.
@@ III.1 Phase diagram
This section presents the phase diagram of our model for $p=1$ (one stored memory). Initially, we analyze the number of fixed points in the time evolution of $M_{Z}$ [Eq. (24)] and $M_{Y}$ [Eq. (25)] as follows:
@@
Additionally, we have $M_{Y}=-4\Omega M_{Z}$. Notably, the origin $(M_{Z},M_{Y})=(0,0)$ is always a fixed point. The boundary at which the number of fixed points changes is determined when the left-hand side and right-hand side of Eq. 28 are tangential:
@@
The parameters $(T,\Omega)$ that satisfy these two equations provide the desired boundary.
@@ III.1.1 $x=2$
This defines the boundary where the number of fixed points varies. When $\beta<\beta_{c}$, the origin is the only fixed point; however, when $\beta>\beta_{c}$, three fixed points exist.
We also numerically examine the region with the LC, as shown in Fig. 1, where the area above the black line corresponds to a region with a single solution, while the area below has two or more solutions. However, after examining the stability of each solution, we identify a region below the black line where the ferromagnetic (FM) solution is unstable. Therefore, we perform a simulation and obtain the green and blue regions, as shown in Fig. 1. The green region in Fig. 1 represents the LC phase, while the blue region is the FM+LC phase. Furthermore, the boundary obtained by calculating the stability of the FM solutions in Ref. [21] does not coincide with the boundary between the blue and green regions in Fig. 1.
Table 1: Comparison of the phase at $x=2$ and $x=4$.
Number $x=2$ [21] $x=4$ (this paper)
$(1)$ FM FM
$(2)$ LC PM + LC
$(3)$ FM + LC FM + LC / PM + LC
$(4)$ PM PM
@@
Consequently, the stability of the PM solution is given by the following conditions [21]:
@@
    $\{\beta<\beta_{c}\}\cap\{\beta>3/2\}\cap\{\beta>4\Omega+1/2\}$: Since the eigenvalues are positive real values, the solution is unstable.
@@
Conditions 1 and 2 indicate that the PM solution is unstable and are classified as the FM phase. Condition 3 indicates instability with spiraling behavior and is classified as the LC phase. Conditions 4 and 5 indicate stability and are classified as the PM phase.
Figure 1: Phase diagram of the quantum modern discrete Hopfield network ($x=2$) in $(T,\Omega)$ plane. The black line indicates the boundary [Eq. (31)] where the number of solutions changes. The blue+green region signifies the area where limit cycles (LCs) are detected from an initial value away from the origin ($(M_{Z},M_{Y})=(3,-3)$). The green region is the area where LCs are detected from an initial value around the origin ($(M_{Z},M_{Y})=(0.05,-0.05)$). Alternatively, the blue region is where the initial values outside the LC converges to it, while the initial value inside the LC and close to the origin converges to the stable fixed points, indicating the ferromagnetic + LC area. The numbering from $(1)$ to $(4)$ corresponds to Table 1.
@@ III.1.2 $x=4$
which determines the boundary at which the number of fixed points changes. When $\beta>\beta_{c}$, there are five fixed points instead of three owing to the nontrivial power inside $\tanh(\cdot)$.
@@
Figure 2: Phase diagram of the quantum modern discrete Hopfield network ($x=4$) in $(T,\Omega)$ plane. The black line indicates the boundary [Eq. 36], where the number of solutions changes. The blue region is the area where the simulation suggests the presence of the limit cycle (LC) when the initial value is $(M_{Z},M_{Y})=(3,-3)$. Only one stable attractor in the paramagnetic (PM) phase (above the curve) and the ferromagnetic phase has two or more stable attractors (below the curve). The label LC + PM indicates the region where the LC appears and the origin remains stable.
Figure 3: Enlarged view of the region where the limit cycle (LC) appears. $(1)$ to $(4)$ represent the number in Table 1, $(2)$ refers the blue region above the black line and $(3)$ refers to the narrow blue region below the black line. The region $(3)$ contains two different phases, depending on whether the fixed points with $M_{Z}>0$ are stable (ferromagnetic+LC) or unstable (paramagnetic+LC).
@@
This transformation allows us to obtain the stability around the solution, classifying it similarly to the case for $x=2$. However, this boundary changes slightly from the existing boundary.
In addition to the fixed points, we numerically identify some parameter regions with a nontrivial LC encircling the origin (blue region in Figs. 2 and 3). However, unlike in the open quantum Hopfield network, the LC consistently coexists with at least one stable fixed point (origin), and the shape of the region in the phase space differs qualitatively from that of the quantum Hopfield network.
@@ III.2.1 PM phase
The PM phase corresponds to the region where every initial state reaches the origin, which is the only stable fixed point ((d) and (e) in Fig. 4).
III.2.2 PM+LC phase
@@
The PM+LC phase corresponds to the region where the only stable fixed point is the origin; however, there is an additional LC around it ((a), (b) and (c) in Fig. 5). In a quantum Hopfield network, the LC phase does not coexist with the PM phase; thus, any initial $(M_{Z},M_{Y})$ converges to the LC. However, in our model, the LC always coexists with other stable fixed points. In the PM+LC phase, the initial point converges to the origin or the LC depending on its proximity to the origin.
III.2.3 FM phase
@@
The region with stable fixed points, such that $M_{Z}>0$, but without the LC, is termed the FM phase ((a), (b) and (c) in Fig. 4). Unlike the FM phase of a quantum Hopfield network [21], which only has stable fixed points with $M_{Z}>0$, our model exhibits the FM phase with an additional stable fixed point at the origin. Consequently, if the initial $M_{Z}$ is sufficiently small, the magnetization $M_{Z}$ diminishes, similar to the PM phase. In addition, the FM region in Fig. 2 is smaller than that in a quantum Hopfield network [21].
III.2.4 FM+LC phase
@@
The FM+LC phase corresponds to the region where FM and LC coexist ((d) in Fig. 5). In this region, LC encapsulates all stable fixed points. When the initial point lies within the LC, it converges to one of the fixed points or the LC, depending on the closeness to these points. Conversely, any point outside the LC converges to the LC.
III.2.5 Behavior of the fixed points for larger $x$
@@
We also analyze the time evolution for $x$ values exceeding 44Fig. 6). In a discrete modern Hopfield network, memory capacity increases nonlinearly with $N$ as $x$ increases [7]. The time evolution of the averaged spins varies with increasing $x$ in an open quantum discrete Hopfield network. The point of change is that the range of $M_{Z}$ that converges to the origin expands as the value of $x$ increases.
@@
The open quantum discrete modern Hopfield network is described by Eqs. (26) and (27), and the behavior of these functions changes only with $x$ via $\tanh(\beta(M_{Z})^{x-1})$. The function $\tanh(\beta(M_{Z})^{x-1})$ has no extreme value at $x=2$, but has an extreme value $M_{Z}=0$ when $x>2$, indicating that the origin is always stable. Consequently, the behavior of the model changes qualitatively when moving from $x=2$ to $x>2$, but remains qualitatively unchanged for different values of $x>2$.
@@
Figure 4: The time evolution of the averaged spin for each parameter $T,\Omega$ in the ferromagnetic (FM) and paramagnetic (PM) phases. The red dots are the stable fixed points. (a), (b) and (c) represent the time evolution in the FM phase, and (d) and (e) represent the time evolution in the PM phase.
@@
Figure 5: The time evolution of the averaged spin for each parameter $T,\Omega$ in the phase in which the limit cycle (LC) appears. (a), (b) and (c) represent the time evolution of the averaged spin in the paramagnetic+LC phase, and (d) represents the time evolution of the averaged spin in the ferromagnetic+LC phase.
@@ III.3 Storage capacity
In the problem of associative memory, storage capacity is the maximum number of patterns that can be retrieved. When considering a landscape of energy function, each stored pattern is an attractor of that landscape. Two methods are used to analyze the storage capacity of Hopfield networks. The first calculates the maximum number of patterns that can be stored without unacceptable errors [1, 28]. The second employs statistical physics analysis [5, 28]. Notably, the Hopfield network can store $0.14N$ patterns.
The discrete modern Hopfield network can store $\alpha_{x}N^{x-1}$ patterns where $\alpha_{x}$ is a numerical constant that depends on the threshold of error [7]. Moreover, models with more generalized energy functions and time evolution can achieve an exponential storage capacity [8, 9]. The storage capacity $\alpha_{c}$ of a quantum Hopfield network at a given temperature $T$ and quantum drive $\Omega$ is also known [25] (note that $\alpha_{c}$ is the number of patterns per node, which is sometimes called the storage capacity). The storage capacity of our model can be calculated using the method described in Ref. [25]. While the classical and quantum limits coincide, further exploration of quantum capacity remains a future challenge.